how do solve and graph these answers on the number line..m+1 3/8=5 what is can you show me step by step how to solve this problem.
Answer by stanbon(75887)

How do you solve and graph these answers on the number line..
m + 1 3/8=5
Can you show me step by step how to solve this problem.
------------------
m + (11/8) = (40/8)
Subtract 11/8 from both sides to get:
---
m = 29/8
m = 3 5/8
=========================================
Cheers,
Stan H.